OpenStudy (anonymous):

F(x)={x3 x<=2 , ax2 x>2 F(x)={2 x<=-1 ,ax+b -1=3 find a,b

OpenStudy (anonymous):

The question is confusion. Is it piece wise function with 3 intervals?

OpenStudy (anonymous):

yeah it is. sorry I don't know how to write it. I do my best

OpenStudy (anonymous):

so, x= 1) 2, x<=-1 2) x^3, x<=2 3) ax^2, x>2 and ax+b-1=3 ?

OpenStudy (anonymous):

1)x^3, x<=2 2)ax^2 , x>2 1)2, x<=-1 2)ax+b , -1<x<3 3)-2, x=>3 a,b=? I think this is better now

OpenStudy (anonymous):

so, there are 2 different problems?

myininaya (myininaya):

i don't think you ever said what you wanted to do with these functions

OpenStudy (anonymous):

there are two function that when we find a then we should find b.

OpenStudy (anonymous):

I want find a & b

myininaya (myininaya):

a and b can be anything

OpenStudy (anonymous):

ok.thanks.maybe I should look for the right question

OpenStudy (anonymous):

makes no sense to me. i wrote it on a piece of paper and i get gibberish. i think the person who asked to question might have been confused

myininaya (myininaya):

yes it seems that way i just thought you could use your esp skills

OpenStudy (anonymous):

maybe it is two different questions and maybe there are instructions we do not see

OpenStudy (anonymous):

for example second one could be F(x) = 2 IF x < -1 ax + b IF -1<x<3 the line -1 = 3 makes pretty much no sense

OpenStudy (anonymous):

but i have one for you i will post it

OpenStudy (anonymous):

It is 2 different function

OpenStudy (anonymous):

question 45 & 46 please